Given the population of 5000 scores with a mean of 86 and standard deviation of 10. How many scores


1
Expert's answer
2021-10-21T13:29:10-0400

a)


"P(86<X<96)=P(Z<\dfrac{96-86}{10})-P(Z\leq\dfrac{86-86}{10})"

"=P(Z<1)-0.5=0.341345"

"0.341345(5000)=1706.7\approx1707"

b)


"P(96<X<106)=P(Z<\dfrac{106-86}{10})-P(Z\leq\dfrac{96-86}{10})"

"=P(Z<2)-P(Z\leq 1)=0.977250-0.841345"

"=0.135905"

"0.135905(5000)=679.5\approx680"

c)


"0.56(5000)=2800"

2800 scores are below 56%

"P(X<56)=P(Z<\dfrac{56-86}{10})=P(Z<-3)"

"=0.001350"

"0.001350(5000)=6.75\approx7"

7 scores are below 56.